OpenStudy (howard-wolowitz):

first work out the mean value of the lenghts then work out the Standard Deviation By the empirical rule 99.7% of the population of alligators should lie between the mean MINUS 3*SD and the mean PLUS 3*SD

OpenStudy (howard-wolowitz):

heres my question: Consider the lengths of alligators to be { 18.7, 12.3, 18.6, 16.4, 15.7, 18.3, 15.8, 14.9, 17.6, 16.4, 16.7, 17.8, 16.2, 13.8, 15.2, 14.7, 13.2, 15.8, 14.6, 16.6, 16.2, 18.1, 17.5, 17.3}. According to the Empirical Rule, 99.7% of the data should fall between what two numbers? A. 11.1 and 21.3 B. 12.3 and 18.1 C. 12.8 and 19.6 D. 13.4 and 18.8

OpenStudy (howard-wolowitz):

now i do not know how to find the value length

hartnn (hartnn):

mean = 16.18, SD = 1.69 mean + 3SD = ... mean - 3SD = ... so what exactly did you not know?

OpenStudy (howard-wolowitz):

someone tld me to find the value lengths

hartnn (hartnn):

mean + 3SD = 16.18 + 3*1.69 = 21.25 mean - 3SD = 16.18 - 3*1.69 = 11.11 so A. wasn't difficult or confusing :P

OpenStudy (howard-wolowitz):

i keep getting high numbers doing that

OpenStudy (howard-wolowitz):

im gonna do the next one by myself and see if i can get a answer